OpenStudy (anonymous):

Which of the following is equivalent to log 4 32 ? Answer A)8 B)2.5 C)0.4 D)1.726

OpenStudy (anonymous):

\[\log_4(16)=x\iff 4^x=16\implies x=2\]

OpenStudy (anonymous):

oops you had 32!!

Directrix (directrix):

B)2.5 4^x = 32 (2^2) ^x = 2^5 2^(2x) = 2^5 2x = 5 x = 5/2 = 2.5

OpenStudy (anonymous):

\[\log_4(32)=x\iff 4^x=32\implies 2^{2x}=32\implies 2x=5\implies x=\frac{5}{2}\]

OpenStudy (anonymous):

Thank you guys!

Directrix (directrix):

We're glad to help.